OpenStudy (sanjanap):

Stats Help! Write 5 positive numbers with a mean of 7 and median of 10

OpenStudy (amistre64):

well, what have you tried?

OpenStudy (amistre64):

since we dont need different numbers, just a set of them that fit the bill x,x,10,y,y may work

OpenStudy (sanjanap):

I cam e up withe some numbers but...they obviously don't work

OpenStudy (sanjanap):

wait nevermind i get it.

OpenStudy (sanjanap):

Thanks you gave me the idea

OpenStudy (amistre64):

mean is (2x + 2y + 10)/5 = 7 2x + 2y + 10 = 35 2x + 2y = 25 x + y = 12.5

OpenStudy (amistre64):

good luck :)